# Integration vs segregation: evidence from measuring and comparing interdisciplinarity in funded and unfunded research using a network approach

### v2

@@ Abstract
Interdisciplinary research fuels innovation. However, research investment has been found to have become more conservative and highly interdisciplinary ideas tend to be punished during research evaluation. How does this affect the interdisciplinarity of research output? To answer this question, our study presents a novel network approach by comparing the evolution of interdisciplinarity in funded research against unfunded research. Considering the field of infectious disease research, the analysis of temporal networks reveals (i) the growing compartmentalisation of funded research, i.e., it becomes more conservative and focuses on the groups of knowledge with readily established connections, and (ii) the growth in global integration in unfunded research, i.e., it tends to be more widely exploratory and links distant knowledge. In addition, we find that in general funded research is less interdisciplinary than unfunded research but this effect decreases through time. Moreover, we observe that interdisciplinary research on prominent infectious diseases like HIV tends to have strong bridging effects facilitating global knowledge integration in the network. In terms of the coronavirus research, despite the surge in publications since COVID-19, we find its systematic impact on knowledge integration remains relatively low. Overall, this research provides detailed insights into the impact of conservatism in funding interdisciplinary research, allows for the identification of under/over-funded interdisciplinary research areas, and offers valuable implications for policymakers aiming to foster interdisciplinary research.
@@
Interdisciplinarity , Knowledge Integration , Funding , Integration , Segregation , Temporal Network
@@
We propose a novel framework to extract and compare funded and unfunded interdisciplinary research in any field of interest and identify potentially under/over-funded interdisciplinary research areas.
We show that conservatism in research funding leads to a compartmentalised structure of funded interdisciplinary research, i.e., established clusters of knowledge get reinforced. On the other hand, unfunded research is generally more exploratory and links more distant fields of study.
We examine how important real-world events influence interdisciplinary research, i.e., we study how interdisciplinary research efforts react to the 2002-2004 SARS outbreak, the 2015-2016 zika outbreak and the 2019-2022 COVID-19 pandemic.
@@ 1 Introduction
Interdisciplinary research (IDR), a process of knowledge integration (Rafols and Meyer, 2010), has been seen as a source of creativity and innovativeness (Rousseau et al., 2019). Various policy and funding initiatives have been developed to facilitate IDR (Wang and Shapira, 2015). However, it has been discovered that highly interdisciplinary (Bromham et al., 2016) and highly novel (Boudreau et al., 2016) ideas that integrate knowledge in unprecedented ways (Uzzi et al., 2013) tend to be penalised during grant proposal evaluation. Given such a lack of appreciation towards highly interdisciplinary ideas at the stage of funding application, does it carry further to research output generated by funding? It is an important question to ask as proposals represent the intention of a scientific idea while publications represent the realisation (Packalen and Bhattacharya, 2020), i.e., when the intention of knowledge integration is discouraged, what happens to the realisation of knowledge integration?
Funding’s influences have been explored in a wide range of aspects, including on research methods and goals (Serrano Velarde, 2018; Luukkonen and Thomas, 2016), team composition (Davies et al., 2022), scientific productivity (Hottenrott and Lawson, 2017), and scholarly, technological and social impact of the associated research output (Heyard and Hottenrott, 2021; Yang, 2024). Of those studies examining various aspects related to research output produced by funding (Jacob and Lefgren, 2011; Wahls, 2019; Packalen and Bhattacharya, 2020; Arora and Gambardella, 2006; Heyard and Hottenrott, 2021; Arora et al., 2000; Benavente et al., 2012; Yang, 2024), limited attention has been paid to examine its interdisciplinarity (ID). In addition, studying funded research output beyond individual funder level at scale remains a gap (Yang, 2024).
In this paper, we address the above gaps. By identifying all funded and unfunded publications in the field of infectious disease from 1995 to 2022, we propose a novel temporal network approach to characterise and compare the ID of funded and unfunded research. We aim to answer the following questions:
  1. 1.
    Are funded research less interdisciplinary in general compared to unfunded research? What are the most under/over-funded interdisciplinary research areas?
  2. 2.
    What are the main trends and tendencies in the time-dependence of ID in funded and unfunded research?
  3. 3.
    Can we identify different temporal regimes in the evolution of ID in funded and unfunded research?
  4. 4.
@@ 2.1 Interdisciplinary Research: Overview
IDR, a process of knowledge integration, generates creativity and innovativeness (National Academy of Sciences et al., 2005; Rafols and Meyer, 2010; Rousseau et al., 2019). Current research on IDR has concluded that IDR could take place cognitively or socially (Glänzel and Debackere, 2022), meaning the integration either happens in a scientist’s mind (Rafols and Meyer, 2010) or in a social process (Abramo et al., 2012, 2017), i.e., formation of a team of researchers with different expertise, experiences, or academic background. Based on these two perspectives, researchers have been exploring a broad range of questions on IDR: from the quantification of the ID of articles (Rafols and Meyer, 2010), journals (Leydesdorff, 2007; Rodríguez, 2017; Leydesdorff et al., 2018), scientists (Porter et al., 2007), institutes (Rafols et al., 2012; Soós and Kampis, 2012), grant proposals (Bromham et al., 2016; Nichols, 2014), and research fields (Leydesdorff and Rafols, 2011), to mapping the global structure of science (Leydesdorff and Rafols, 2009; Rafols et al., 2010), further to the scholarly and social impact of IDR (Okamura, 2019; Hu et al., 2024; Shi et al., 2009; Wang et al., 2015; Yegros-Yegros et al., 2015). For example, Rafols and Meyer (2010) used diversity and coherence as a framework to compare the ID of a set of bionanoscience articles. Leydesdorff and Rafols (2009) and Leydesdorff (2007) analysed the ID of journals based on betweenness centrality and diversity. Porter et al. (2007) proposed the measures of integration, reach and specialisation and used them to measure the ID of $47$ researchers. Soós and Kampis (2012) analysed and compared the disciplinary structure of a sample of Hungarian Research Institutions based on the science overlay maps (Rafols et al., 2010).
@@
IDR may lack institutional appreciation due to mono-disciplinary academic structures (Rousseau et al., 2019) and is poorly rewarded by funders (Bromham et al., 2016) due to disciplinary-based evaluations (Rousseau et al., 2019; Woelert and Millar, 2013). Rafols et al. (2012) compared the ID of Innovation Studies (IS) units with leading Business & Management Schools (BMS) in the UK. Despite IS units being more interdisciplinary, they were disadvantaged in evaluation and obtaining resources due to the bias of the Association of Business Schools’ (ABS) journal rankings favouring mono-disciplinary research. IDR has been found to exhibit higher scholarly impact (Okamura, 2019) but Shi et al. (2009), Wang et al. (2015) and Yegros-Yegros et al. (2015) reported mixed results across different measures (Wang et al., 2015; Yegros-Yegros et al., 2015) and different fields (Shi et al., 2009). IDR has also been found to attract more attention from policy documents (Hu et al., 2024).
@@ 2.3 Funded research and interdisciplinarity
The influence of funding on research has been studied in various aspects (see Thelwall et al. (2023) for further detail), including research methods or goals (Serrano Velarde, 2018; Luukkonen and Thomas, 2016), team composition (Davies et al., 2022), researcher, team and institutional level productivity (Heyard and Hottenrott, 2021; Hottenrott and Lawson, 2017), output types (Thelwall et al., 2023), and scholarly, technological and social impacts of outputs (Heyard and Hottenrott, 2021; Yang, 2024).
@@ 3 Data and Methods
3.1 Data source and curation
@@
Infectious diseases continue to be a major global health issue with significant impact on human society (Baker et al., 2022). Following one of our co-author’s work Head et al. (2020), we based our study on research into 34 major infectious diseases. [^2] The authors of Head et al. (2020) selected a group of 37 prominent infectious diseases for their funding versus disease burden analysis. Our selection excluded sexually transmitted infections and enteric infections because they are a higher level concept than others. We also added consideration of causal relations: trachoma is caused by chlamydia trachomatis but chlamydia also happens to be in the set, so we excluded trachoma. In addition here, we included coronavirus and diphtheria. The inclusion of coronavirus is due to its recent prominent role in the field of infectious disease research. The inclusion of diphtheria is because two other highly related terms, pertussis and tetanus, have already been included in the group (the DTP combination vaccine prevents against all three diseases and they are thus commonly considered together; the three diseases will be addressed by DTP onwards). This gives us 36 selected infectious diseases, which account for \$70 billion infectious disease research funding (65% of overall total) from G20 countries between 2000 and 2017 (Head et al., 2020), and 94% of the total infectious disease burden [^3] Measured in Disability-Adjusted Life Years (DALYs). One DALY represents the loss of the equivalent of one year of full health. in the 2021 Global Burden of Disease data [^4] Available from https://vizhub.healthdata.org/gbd-results/. (Date last accessed: 2024.9.4).
@@
We selected the Web of Science (WoS) Core Collection database as our data source due to its extensive usage in the scientometrics studies (Birkle et al., 2020). We extracted, for each of the 36 diseases, both funded research and all research, the number of records containing the disease in their title, abstract, or author keywords, i.e., the number of occurrences. We also extracted, for both funded and all research, the number of records containing each possible pair of diseases in their title, abstract, or author keywords, i.e., the number of co-occurrences. To retrieve maximum funded records efficiently, we adopt the method of the right-hand truncation search strategy in both the funding agency (FO) and funding grants (FG) field suggested in Liu et al. (2020), as detailed in Search Query 1. We refer to papers that have FO or FG information in the WoS as funded research (labelled “F”) [^5] All research will be labelled “A” hereafter. and those that don’t as unfunded research (labelled “U”). We searched only the document types of original research article, data paper, and proceeding paper, and we specifically excluded any withdrawn or retracted publications, as detailed in Search Query 2. The timespan considered in this paper is from 1995 to 2022 inclusive [^6] The WoS did not include the author keywords for articles until 1991. The years 1991-1994 were excluded from our analysis as the link profiles of these years were very volatile and did not seem to form coherent clusters with any of the rest of the years like the ones in Figure 1.. See Table Data availability for the search strategy for each disease [^7] Search terms for each disease are firstly picked based on experts’ advice and the entry terms (a list of synonyms) in the Medical Subject Heading (MeSH) database. A further selection of search terms was done after testing the added benefit of the search terms, i.e., search terms with low or no additionally identified papers were dropped. [^8] The search term of herpes in our study contains infections related to Herpes Simplex Viruses (HSV) and Shingles. as well as the total and funded number of records returned.
@@
The extracted occurrences and cooccurrences of infectious diseases were recorded in tensors $(C_{F})_{36\times 36\times 28}$ for F and $(C_{A})_{36\times 36\times 28}$ for A. An entry $c_{ij}^{t}$ of $C_{F}$ (or $C_{A}$) gives the cooccurrence count between disease $i$ and $j$ in F (or A) in time slice $t$ [^9] The 28 time slices correspond to years from 1995 to 2022. if $i\neq j$, and the occurrence of $i$ otherwise. The tensor of U $(C_{U})_{36\times 36\times 28}$ is a direct result of element-wise subtraction between $C_{A}$ and $C_{F}$. The data extraction was performed using the Web of Science API Lite [^10] The Web of Science API Lite: support search and data integration using Web of Science data returned as JSON or XML. https://developer.clarivate.com/apis/woslite (Date last accessed: 2024.9.4) in Python. For abbreviations of diseases used in this study, please refer to Table Data availability.
We used one of the most widely-used normalisation measures in scientometrics to normalise the cooccurrence matrix, namely the cosine (Eck and Waltman, 2009) (also known as the Ochiai index (Zhou and Leydesdorff, 2016) or Salton’s index (Adnani et al., 2020)), i.e., $w_{ij}^{t}=\frac{c_{ij}^{t}}{\sqrt{c_{ii}^{t}c_{jj}^{t}}}$. In our setting, the cosine normalises the cooccurrences between $i$ and $j$ at time slice $t$ by the geometric mean of the occurrences of $i$ and $j$ at $t$ when $i\neq j$, and thus measures the correlation between numbers of publications about diseases. When $i=j$, $w_{ij}^{t}=0$. By normalising $C_{A}$, $C_{F}$ and $C_{U}$, we ended up with $W_{A}$, $W_{F}$, $W_{U}$, based on which we then performed the following analysis. Note that we regarded the collection of all WoS publications related to infectious diseases as the system of infectious disease knowledge, where $w_{ij}^{t}$ is the association, or the cognitive proximity, or the current extent of knowledge integration between disease $i$ and $j$ at time $t$.

### v3

# Integration vs segregation: network analysis of interdisciplinarity in funded and unfunded research on infectious diseases

@@ Abstract
Interdisciplinary research fuels innovation. In this paper, we examine the interdisciplinarity of research output driven by funding. Considering 36 major infectious diseases, we model interdisciplinarity through temporal correlation networks based on funded and unfunded research from 1995-2022. Using hierarchical clustering, we identify coherent periods of time or regimes characterised by important research topics like vaccinations or the Zika outbreak. We establish that funded research is less interdisciplinary than unfunded research, but the effect has decreased markedly over time. In terms of network growth, we find a tendency of funded research to focus on readily established connections leading to compartmentalisation and conservatism. In contrast, unfunded research tends to be exploratory and bridge distant knowledge leading to knowledge integration. Our results show that interdisciplinary research on prominent infectious diseases like HIV and tuberculosis tends to have strong bridging effects facilitating global knowledge integration in the network. At the periphery of the network, we observe the emergence of vaccination-related and Zika-related knowledge clusters, both with limited systemic impact. We further show that despite the surge in publications related to COVID-19, its systematic impact on the disease network remains relatively low. Overall, this research provides a generalisable framework to examine the impact of funding in interdisciplinary knowledge creation. It can assist in priority setting, for example with horizon scanning for new and emerging threats to health, such as pandemic planning. Policymakers, funding agencies, and research institutions should consider revamping evaluation systems to reward interdisciplinary work and implement mechanisms that promote and support intelligent risk-taking.
@@
Interdisciplinary Research , Knowledge Integration , Research funding , Conservatism , Temporal Network Analysis
@@
We present a framework to analyse funded vs unfunded interdisciplinary research (IDR)
Funded IDR compartmentalises by reinforcing existing research clusters.
Unfunded IDR is more exploratory by linking distant knowledge.
As a result, funded IDR segregates while unfunded IDR integrates.
We explore how major events like SARS, Zika, and COVID-19 shape IDR.
@@ 1 Introduction
Interdisciplinary research (IDR), a process of knowledge integration (Rafols and Meyer, 2010), has been seen as a source of creativity and innovativeness (Rousseau et al., 2019). Various policy and funding initiatives have been developed to facilitate IDR (Wang and Shapira, 2015). However, it has been discovered that highly interdisciplinary (Bromham et al., 2016) and highly novel (Boudreau et al., 2016) ideas that integrate knowledge in unprecedented ways (Uzzi et al., 2013; Fontana et al., 2020) tend to be penalised during grant proposal evaluation. Given such conservatism towards highly interdisciplinary ideas at the stage of funding application, does it carry further to research output generated by funding? It is an important question to ask as proposals represent the intention of a scientific idea while publications represent the realisation (Packalen and Bhattacharya, 2020), i.e., when the intention of knowledge integration is discouraged, what happens to the realisation of knowledge integration?
Funding’s influences have been explored in a wide range of aspects, including on research methods and goals (Serrano Velarde, 2018; Luukkonen and Thomas, 2016), team composition (Davies et al., 2022), scientific productivity (Hottenrott and Lawson, 2017), and scholarly (Roshani et al., 2021; Mosleh et al., 2022; Coccia and Roshani, 2024a, b), technological and social impact (Heyard and Hottenrott, 2021; Yang, 2024) of the associated research output. Funded research has been found to exhibit higher scholarly impact than its unfunded counterpart in various fields (Roshani et al., 2021; Mosleh et al., 2022; Coccia and Roshani, 2024a, b).
Of those studies examining various aspects related to research output produced by funding (Jacob and Lefgren, 2011; Wahls, 2019; Packalen and Bhattacharya, 2020; Arora and Gambardella, 2006; Heyard and Hottenrott, 2021; Arora et al., 2000; Benavente et al., 2012; Yang, 2024; Roshani et al., 2021; Mosleh et al., 2022; Coccia and Roshani, 2024a, b), limited attention has been paid to examine its interdisciplinarity (ID). In addition, studying funded research output beyond individual funder level at scale remains a gap (Yang, 2024).
In this paper, we address the above gaps. By identifying all funded and unfunded publications in the field of infectious disease from 1995 to 2022 and considering a set of major infectious diseases to be the unit of analysis, we propose a novel temporal network approach to characterise and compare the ID of funded and unfunded research. We aim to answer the following research questions (RQ)s:
  1. RQ1:
    Can we identify coherent periods of time, i.e., temporal regimes, in the evolution of ID in funded and unfunded research?
  2. RQ2:
    What are the main trends in the time-dependence of ID in funded and unfunded research?
  3. RQ3:
    Since IDR is discouraged at grant application, is funded research less interdisciplinary in general compared to unfunded research? What are the most under/over-funded interdisciplinary research areas?
  4. RQ4:
@@ 2.1 Interdisciplinary Research: Overview
According to the National Academies of Sciences of the USA National Academy of Sciences et al. (2005), IDR is ”a mode of research by teams or individuals that integrates information, data, techniques, tools, perspectives, concepts and/or theories from two or more disciplines or bodies of specialized knowledge to advance fundamental understanding or to solve problems whose solutions are beyond the scope of a single discipline or area of research practice.” IDR, as a process of knowledge integration, generates creativity and innovativeness (National Academy of Sciences et al., 2005; Rafols and Meyer, 2010; Rousseau et al., 2019). Current research on IDR has concluded that IDR could take place cognitively or socially (Glänzel and Debackere, 2022), meaning the integration either happens in a scientist’s mind (Rafols and Meyer, 2010) or in a social process (Abramo et al., 2012, 2017), i.e., formation of a team of researchers with different expertise, experiences, or academic background. Based on these two perspectives, researchers have been exploring a broad range of questions on IDR: from the quantification of the ID of articles (Rafols and Meyer, 2010), journals (Leydesdorff, 2007; Rodríguez, 2017; Leydesdorff et al., 2018), scientists (Porter et al., 2007; Leahey et al., 2017), institutes (Rafols et al., 2012; Soós and Kampis, 2012; Biancani et al., 2018), grant proposals (Bromham et al., 2016; Nichols, 2014), and research fields (Leydesdorff and Rafols, 2011), to mapping the global structure of science (Leydesdorff and Rafols, 2009; Rafols et al., 2010), further to the scholarly and social impact of IDR (Okamura, 2019; Hu et al., 2024; Shi et al., 2009; Wang et al., 2015; Yegros-Yegros et al., 2015; Biancani et al., 2018; Leahey et al., 2017). For example, Rafols and Meyer (2010) used diversity and coherence as a framework to compare the ID of a set of bionanoscience articles. Leydesdorff and Rafols (2009) and Leydesdorff (2007) analysed the ID of journals based on betweenness centrality and diversity. Porter et al. (2007) proposed the measures of integration, reach and specialisation and used them to measure the ID of $47$ researchers. Soós and Kampis (2012) analysed and compared the disciplinary structure of a sample of Hungarian Research Institutions based on the science overlay maps (Rafols et al., 2010).
@@
IDR may lack institutional appreciation due to mono-disciplinary academic structures (Rousseau et al., 2019) and is poorly rewarded by funders (Bromham et al., 2016) due to disciplinary-based evaluations (Rousseau et al., 2019; Woelert and Millar, 2013; Rylance, 2015; Fontana et al., 2022). Rafols et al. (2012) compared the ID of Innovation Studies (IS) units with leading Business & Management Schools (BMS) in the UK. Despite IS units being more interdisciplinary, they were disadvantaged in evaluation and obtaining resources due to the bias of the Association of Business Schools’ (ABS) journal rankings favouring mono-disciplinary research. Biancani et al. (2018) found interdisciplinary research centres have superior performance in knowledge production, instruction, collaboration and funding acquisition. Leahey et al. (2017) found that being interdisciplinary is a high risk high reward endeavour in a scientific career, i.e., a trade-off between productivity and scientific impact. IDR has been found to exhibit higher scholarly impact (Okamura, 2019) but Shi et al. (2009), Wang et al. (2015) and Yegros-Yegros et al. (2015) reported mixed results across different measures (Wang et al., 2015; Yegros-Yegros et al., 2015) and different fields (Shi et al., 2009). IDR has also been found to attract more attention from policy documents (Hu et al., 2024).
@@ 2.3 Funded research and interdisciplinarity
The influence of funding on research has been studied in various aspects (see Thelwall et al. (2023) for further detail), including research methods or goals (Serrano Velarde, 2018; Luukkonen and Thomas, 2016), team composition (Davies et al., 2022), researcher, team and institutional level productivity (Heyard and Hottenrott, 2021; Hottenrott and Lawson, 2017), output types (Thelwall et al., 2023), and scholarly (Roshani et al., 2021; Mosleh et al., 2022; Coccia and Roshani, 2024a, b), technological and social impacts of outputs (Heyard and Hottenrott, 2021; Yang, 2024).
Funding sources, especially those that routinely issue topic-focused calls (Viergever and Hendriks, 2016), seem to have the greatest influence on most science and health research which are resource intensive (Thelwall et al., 2023; Whitley et al., 2018); however, the degree of such influence depends on individual funders’ requirements (Serrano Velarde, 2018; Luukkonen and Thomas, 2016; Thelwall et al., 2023).
Funding has demonstrated great value to knowledge creation by facilitating scientific productivity (Heyard and Hottenrott, 2021; Hottenrott and Lawson, 2017). Funding also plays a key role in knowledge diffusion (Roshani et al., 2021; Mosleh et al., 2022; Coccia and Roshani, 2024a, b), where funded research tends to be more often cited than unfunded research in a wide variety of fields including computer science, economics, physics, chemistry and medicine. Moreover, funded research also tends to produce higher technological impact (through citation by patent) and social engagement (through citation by Tweets) (Yang, 2024).
@@ 3 Data and Methods
3.1 Data source and extraction procedure
@@
Infectious diseases continue to be a major global health issue with significant impact on human society (Baker et al., 2022), for example the COVID-19 pandemic, and the threat of climate change upon human health and infections (Mora et al., 2022). IDR is required to address this multifaceted challenge (Wilcox and Colwell, 2005). Knowledge integration across infectious diseases has already contributed greatly to clinical and research knowledge (Schwetz and Fauci, 2019). Following one of our co-author’s work Head et al. (2020), we based our study on research into 34 major infectious diseases. [^2] The authors of Head et al. (2020) selected a group of 37 prominent infectious diseases for their funding versus disease burden analysis. Our selection excluded sexually transmitted infections and enteric infections because they are a higher level concept than others. We also added consideration of causal relations: trachoma is caused by chlamydia trachomatis but chlamydia also happens to be in the set, so we excluded trachoma. In addition here, we included coronavirus and diphtheria. The inclusion of coronavirus is due to its recent prominent role in the field of infectious disease research. The inclusion of diphtheria is because two other highly related terms, pertussis and tetanus, have already been included in the group (the DTP combination vaccine prevents against all three diseases and they are thus commonly considered together; the three diseases will be addressed by DTP onwards). This gives us 36 selected infectious diseases which were considered as the units of analysis in this paper. They account for \$70 billion infectious disease research funding (65% of overall total) from G20 countries between 2000 and 2017 (Head et al., 2020), and 94% of the total infectious disease burden [^3] Measured in Disability-Adjusted Life Years (DALYs). One DALY represents the loss of the equivalent of one year of full health. in the 2021 Global Burden of Disease data [^4] Available from https://vizhub.healthdata.org/gbd-results/. (Date last accessed: 2024.9.4).
We selected the Web of Science (WoS) Core Collection database as our data source due to its extensive usage in the scientometrics studies (Birkle et al., 2020). The collection of publications on each disease can be seen as a body of knowledge (Hamburg, 2008), and thus the co-occurrence of diseases in the title, abstract or author keywords represents the practice of knowledge integration (National Academy of Sciences et al., 2005). We extracted, for each of the 36 diseases, both funded research and all research, the number of records containing the disease in their title, abstract, or author keywords, i.e., the number of occurrences. We also extracted, for both funded and all research, the number of records containing each possible pair of diseases in their title, abstract, or author keywords, i.e., the number of co-occurrences.
@@
To retrieve maximum funded records efficiently, we adopt the method of the right-hand truncation search strategy in both the funding agency (FO) and funding grants (FG) field suggested in Liu et al. (2020), as detailed in Search Query 1. We refer to papers that have FO or FG information in the WoS as funded research (labelled “F”) [^5] All research will be labelled “A” hereafter. and those that don’t as unfunded research (labelled “U”). We searched only the document types of original research article, data paper, and proceeding paper, and we specifically excluded any withdrawn or retracted publications, as detailed in Search Query 2. The timespan considered in this paper is from 1995 to 2022 inclusive [^6] The WoS did not include the author keywords for articles until 1991. The years 1991-1994 were excluded from our analysis as the link profiles of these years were very volatile and did not seem to form coherent clusters with any of the rest of the years like the ones in Figure 1.. See Table Data availability for the search strategy for each disease [^7] Search terms for each disease are firstly picked based on experts’ advice and the entry terms (a list of synonyms) in the Medical Subject Heading (MeSH) database. A further selection of search terms was done after testing the added benefit of the search terms, i.e., search terms with low or no additionally identified papers were dropped. [^8] The search term of herpes in our study contains infections related to Herpes Simplex Viruses (HSV) and Shingles. as well as the total and funded number of records returned.
@@
To explore the possibility that an instance of co-occurrence of diseases in the abstract of a paper is not because they are the primary research objective of the paper but merely a mention as an aside we carried out some manual checks on a sample of papers. For this purpose, we selected five pairs of representative infectious diseases (HIV-TB, Dengue-Zika, Tetanus-Diphtheria, HCV-HBV, Clamydia-Gonorreahea) which have important contributions to the system dynamics and performed a sample test of size 100 on each pair to examine the proportion of wrong identifications. As shown in Table 1, we found that the false positive rates are consistently less than 3% and we found no bias towards particular disease pairs.
The extracted occurrences and cooccurrences of infectious diseases were recorded in tensors $(C_{F})_{36\times 36\times 28}$ for F and $(C_{A})_{36\times 36\times 28}$ for A. An entry $c_{ij}^{t}$ of $C_{F}$ (or $C_{A}$) gives the cooccurrence count between disease $i$ and $j$ in F (or A) in time slice $t$ [^9] The 28 time slices correspond to years from 1995 to 2022. if $i\neq j$, and the occurrence of $i$ otherwise. The tensor of U $(C_{U})_{36\times 36\times 28}$ is a direct result of element-wise subtraction between $C_{A}$ and $C_{F}$. The data extraction was performed using the Web of Science API Lite [^10] The Web of Science API Lite: support search and data integration using Web of Science data returned as JSON or XML. https://developer.clarivate.com/apis/woslite (Date last accessed: 2024.9.4) in Python, and the network analysis was performed using R. For abbreviations of diseases used in this study, please refer to Table 1.
We used one of the most widely-used normalisation measures in scientometrics to normalise the cooccurrence matrix, namely the cosine (Eck and Waltman, 2009) (also known as the Ochiai index (Zhou and Leydesdorff, 2016) or Salton’s index (Adnani et al., 2020)), i.e., $w_{ij}^{t}=\frac{c_{ij}^{t}}{\sqrt{c_{ii}^{t}c_{jj}^{t}}}$. In our setting, the cosine normalises the cooccurrences between $i$ and $j$ at time slice $t$ by the geometric mean of the occurrences of $i$ and $j$ at $t$ when $i\neq j$, and thus measures the correlation between numbers of publications about diseases. When $i=j$, $w_{ij}^{t}=0$. By normalising $C_{A}$, $C_{F}$ and $C_{U}$, we ended up with $W_{A}$, $W_{F}$, $W_{U}$, based on which we then performed the following analysis. Note that we regarded the collection of all WoS publications related to infectious diseases as the system of infectious disease knowledge, where $w_{ij}^{t}$ is the association, or the cognitive proximity, or the current extent of knowledge integration between disease $i$ and $j$ at time $t$. A high value of $w_{ij}^{t}$ implies an established knowledge link between the pair while a low value suggests a non-established state.